# Future Trends in the Design of Memetic Algorithms: the Case of the Linear Ordering Problem

### v1

@@ 4 Adaptations of MA-EDM to Improve its Long-Term Effectiveness
As computing power increases and the number of instances grows, it is quite common for the effectiveness of algorithms to be affected [17]. Different options to deal with the increase in computing power, comparing the alternatives for expanding the number of generations or the size of the population have been analized [46]. Based on its analyses, both theoretical and practical reasons are given in favor of increasing the size of the population and keeping the number of generations from growing excessively. There are currently attempts to develop effective population-based algorithms with large populations, since this would greatly facilitate the efficient use of computational resources, especially in parallel computing environments [47]. However, the use of large populations is still in the minority. Moreover, we know that by introducing mechanisms that delay premature convergence, the number of useful generations can be increased [32]. For example, some optimizers that use Best Non-Penalized (BNP) strategy — such as MA-EDM — have been able to successfully leverage several thousand generations [34]. However, even with this type of mechanism, there are limits on the number of useful generations, so with large computing capacities, some adaptations are still necessary.
The options to take advantage of the increased computing power to manage larger populations or handle more generations do not entail any type of change in the algorithm. A different option that will also be evaluated in this research to avoid the evolution of such a large number of generations is to introduce computationally effective (and costly) strategies in some of the components that are used in each iteration of the algorithm. In particular, for the improvement process, instead of applying a local search whose criterion is based on reaching a local optimum, a trajectory metaheuristic can be used to set a time-based stopping criterion. This option is not promising at present for the LOP with large instances because with the current amount of power, it would require setting a very short time or having very few generations, but as the computing power increases, this option becomes viable. This methodology makes it possible to control the number of generations without having to resort to increasing the size of the population by an inordinate amount.
@@
In order to properly design an ILS strategy, we must select the perturbation and local search components. Given the efficiency and effectiveness of the local search based on shifts, this was left unchanged. In the case of the perturbation strategy, we analyzed several operators that have been successful in combinatorial optimization [49] and that exhibit different perturbation strengths. In particular, we tried strategies widely used in the field of ILS: Scramble, Inversion, Insertion and Swap. In the case of the swap strategy, we considered the version that incorporates the parameter to establish the number of 2-position swaps that are made [50]. This parameter is used to model the strength of the perturbation. The preliminary analyses allowed us to determine that the method based on performing multiple swaps was far superior. In particular, was enough to escape local optima, so the rest of the paper considers this operator exclusively. In fact, this operator had already shown its superiority in [48], but since in this research, the ILS is integrated into the MA, we opted to obtain an independent confirmation.
To extend MA-EDM (Algorithm 1) and generate MA-EDM, we need only change the intensification steps (lines 2 and 7) by replacing the local search with the ILS. The Algorithm 2 details each step of the ILS. Note that it is the standard variant of ILS in which each solution is intensified with a local search after being perturbed, and it is accepted if it is at least as good as the best solution found. In relation to the parallel version of this algorithm, it follows exactly the same diagram presented above (see Figure 1). The only difference is that the workers execute the ILS instead of the local search. The parallel version is called PMA-EDM.
@@ 5 Experimental Validation
The main objective of the experimental validation carried out in this research is to analyze how well the MA-EDM algorithm can be scaled up with an increase in computing power, as well as to understand how its parameterization or its components can be changed to contend with this situation. As a secondary objective, we intend to apply the algorithms developed to the most challenging instances of the LOP by simulating scenarios with a large amount of computation, in order to understand the room for improvement present compared to the current BKS [19], as well as to establish new levels that can be used in the future to validate new advances in this area. The results are presented in two sections, each one focused on each of the objectives listed above.
To account for the stochastic behavior of the algorithms analyzed, the runs were repeated 30 times in all the experiments, and the following statistical analyses were performed to compare the results [51]. First, the Kruskal Wallis test was applied as an omnibus test [52]. If there were any significant differences, the Mann Whitney with Hommel’s correction was used to identify the pair-wise differences. Additionally, with the aim of visualizing the results, the Nemenyi critical difference plot was used [53], with a significance level equal to 0.05. The experimental validation was carried out in a cluster whose nodes have two Intel Xeon E5 v4 family processors and 32 GB of RAM.
@@ 5.1 Relationship between the scalability and components of MA-EDM
In order to analyze the scalability of MA-EDM in anticipation of a future increase in processor computing power, the increase is simulated in two ways: by increasing the stopping criterion, or through parallelization. The former yields an exact analysis of what will happen with an increase in computing power, allowing for a more limited range of increase, while the parallelization simulates a much larger increase in resources, although when there is a loss of efficiency, the value of the increase in computing power is approximate. It should be noted that the parallelization used keeps intact the evolutionary process of the MA that is parallelized.
@@
First, in order to validate the benefit of increasing the number of generations, the performance of the MA-EDM and MA-EDM sequential implementations was analyzed using different stopping criteria (6, 24 and 192 hours). In order to have instances with different complexities, four instances of the xLOLIB2 package were chosen at random, one for each size available: 300, 500, 750 and 1000. The instances used were: N-t65i11xx_300, N-be75eec_500, N-stabu75_750 and N-t70b11xx_1000. Figure 2 shows the boxplots of the results obtained in each case. The only parameter of the MA-EDM algorithm is the population size (). Taking into account the results in [19], was used. In the case of MA-EDM, the same population size was used and the stopping criterion for ILS was set at 3.6 seconds. The preliminary analyses showed that after this run time, the improvements offered by ILS are quite limited. With this value, each generation lasts about 12 minutes, so in the longest runs, about generations evolve. In all the instances, it is evident that in runs with more limited stopping criteria (6 and 24 hours), the MA-EDM algorithm is superior. This is especially clear in the 6-hour run, a case in which there are significant differences in favor of MA-EDM in the 4 instances. However, as the computing power increases, the improvements achieved by the MA-EDM algorithm are quite limited, and it is far surpassed by the MA-EDM algorithm. This improvement is most noticeable for larger instances. In fact, in the instances with , and , the statistical tests mentioned earlier confirm the superiority of MA-EDM.
The advantages offered by MA-EDM in instances of size are more limited. The increase in computing power does not yield an increase in the fitness of the best solutions achieved, but it does produce a noticeable decrease in the standard deviation (both in MA-EDM and MA-EDM), so it is very possible that solutions very close to optimal are being achieved in instances of this size. In the cases of , and , the improvement provided by MA-EDM is broad, with even the worst result achieved by MA-EDM being better than the average achieved by MA-EDM. In the case of the instance of size , MA-EDM and MA-EDM evolved on average and generations, respectively. The results show that as the complexity of the instances grows, expanding the number of generations is not the best way to take advantage of the increased computational resources. In fact, in the case of the instance, despite the noticeable margin for improvement still left for MA-EDM, between the results obtained in 24 hours and 192 hours, i.e., when multiplying the computation amount by 8, there were no significant differences.
@@
Taking into account the analyses carried out in [46], it is natural to consider MAs with large populations as an alternative to leverage the increase in computing power. Because of this, MA-EDM was run by setting the stopping criterion to 192 hours and using larger population sizes. In particular, 6 values were tested between and . The boxplots of the results obtained for the instances N-stabu75_750 and N-t70b11xx_1000 are shown in Figure 3. The results obtained by MA-EDM are also included. The statistical tests carried out confirm that in both instances, there were significant differences between the results achieved by MA-EDM and those obtained by MA-EDM, regardless of the population size used. As expected, there was a decrease in the number of generations. As an example, MA-EDM with averaged generations for instances of size , while increasing to reduced the number of generations to . However, this was not reflected in a so significant improvement in the quality of the solutions. We conclude that in this case, decreasing the number of generations through population growth is not enough to take advantage of the computing resources; instead, more drastic changes must be made by modifying some of its internal components, as is done in MA-EDM.
@@ 5.2 Analysis of performance in the very long-term
The above experiment showed that MA-EDM provides significant improvements in instances of size or larger. This second experiment seeks to establish new BKS and analyze whether MA-EDM scales appropriately when considering larger computation quantities than those used in the previous experiment. Since the previous experiment involved optimizations lasting up to 8 days, it is not feasible to perform this new experiment using sequential executions. Instead, PMA-EDM with 32 worker processes was used. One modification was made to the population size, consisting of establishing in order for the population size to be a multiple of the number of workers. Figure 4 shows the boxplots obtained by PMA-EDM when executed setting the stopping time between 6 hours and 26 hours for the 750 and 1000 size instances that were used in the previous experiment. Also shown are the results obtained by MA-EDM at 192 hours. The results for the 300 and 500 size instances are shown in the Supplementary Material. Note that the computation hours used in the parallel cases range from 192 to 832 hours. As expected, the 6-hour parallel results were worse than the 192-hour sequential results due to the loss of performance inherent in the parallel models. However, when using 10-hour runs, the median obtained by PMA-EDM was higher than that of MA-EDM. Therefore, the acceleration is estimated to be between and . When considering the mean value, the efficiency is equal to 0.8, and therefore each hour of parallel execution is estimated to correspond to hours of sequential execution. Thus, this experiment simulates situations with current computing capabilities between 153 hours and 665 hours, i.e., practically one month of execution.
@@
In the case of the instance of size , the results continue to improve steadily, even on the longest runs. The improvement in quality obtained compared to the previous BKS is immense, thus confirming that in instances of size with the usual execution duration — at most a few days — there is considerable room for improvement in the LOP. Figure 5 shows the plots of the critical difference for these runs for the and instances. These same figures for the case of 500 are shown in the Supplementary Material of this research. In the case of 300, it is not possible to construct this graph since the same result was obtained for all the executions carried out with both algorithms. In both instances, we see that the sequential run at 192 hours ranked between the parallel runs at 6 and 10 hours, confirming the previous findings in terms of the estimated efficiency. In the 750 instance, 22-hour parallel runs were needed to significantly outperform the sequential results. When a method generates near-optimal solutions, the resources allocated have to be greatly increased to achieve significant improvements, so this, coupled with the fact that the 22-hour run ranked similar to the 26-hour run, suggests that in this instance, very high quality solutions are being reached or that the method developed is no longer capable of yielding additional improvements. In the case of the instance, the sequential results at 192 hours could be significantly outperformed by all the parallel runs starting at 14 hours. In addition, the best average ranking was obtained at 26 hours, which is an indication that there is still room for improvement.
| Instance | PMA-EDM | Previous BKS | ||
| Mean | Best | Worst | ||
@@
| Instance | PMA-EDM | BKS | ||
| Mean | Best | Worst | ||
@@
6 Conclusions
The design of metaheuristics has evolved over time due to the emergence of new ideas to perform optimization processes, as well as the continuous increase in computing capabilities. This paper focuses on understanding the impact that the future increase in the availability of computational resources will have on Memetic Algorithms (MAs) in the particular case of the Linear Ordering Problem (LOP). The use of parallelizations that keep the evolutionary model intact, as well as of long-term executions, made it possible to confirm that the MA that is used today to obtain the Best-Known Solutions (BKS) in the most challenging instances of the LOP (MA-EDM) is not able to adequately leverage the increase in computational resources on a large scale. Two of the most common techniques used to take advantage of the increase in computational resources are to expand the size of the population or the number of generations. However, with a high increase in computational resources, neither of these two techniques was successful in the case of the LOP. A new heuristic optimizer (MA-EDM) was proposed based on integrating one of the most successful schemes of the 1990s (Iterated Local Search) with MA-EDM, so as to avoid an excessive increase in the population size or in the number of generations, and thus increase the intensification potential. This technique was able to successfully take advantage of amounts of computational resources similar to what a modern core can execute in more than 4 months, and yielded new solutions that significantly outperform the current BKS in the 18 instances that were used in this research. These achievements confirm that as computational resources continue to increase, it will be necessary to adapt the current optimizers, and in particular it shows how promising it is to integrate MAs with techniques with great intensifying power. Accordingly, we believe that in the future there will be a significant boom in MAs that will include intensification procedures with a high computational cost. In fact, in recent years, the term matheuristics has become popular to refer to heuristics that include mathematical programming techniques in some of their components. These types of techniques were not viable in the past due to their computational cost, but the number of areas in which this type of integration has been successful is growing. Also of note is the fact that the generation of BKS has been possible without incorporating initialization mechanisms or genetic operators that bias the search towards certain types of regions that might seem more promising. These types of ad-hoc components have been successful in the past. However, with the increase in computational resources, more regions can be explored, so the inclusion of some type of bias in the search, specially in the initial phases, can be counterproductive. Therefore, we predict that in the future, operators with a high degree of randomness - such as those used in this paper - will become more effective.

### v2

@@ 4 Adaptations of MA-EDM to Improve its Long-Term Effectiveness
As computing power increases and the instances grow, it is quite common for the effectiveness of algorithms to be affected [9]. Several options to deal with the increase in computing power, including the alternatives for expanding the number of generations or the size of the population have been analyzed [25]. Based on these analyses, both theoretical and practical reasons are given in favor of increasing the size of the population and keeping the number of generations from growing excessively. Alternatively, by introducing mechanisms that delay premature convergence, the number of useful generations can be increased [19]. For example, some optimizers that use Best Non-Penalized (BNP) strategy — such as MA-EDM — have been able to successfully leverage several thousand generations [26]. However, even with this type of mechanism, there are limits on the number of useful generations, so with large computing capacities, some adaptations are still necessary.
The options to take advantage of the increased computing power to manage larger populations or handle more generations do not entail any type of change in the algorithm. A different option is to introduce computationally effective (and costly) strategies in some of the components that are used in each iteration of the algorithm. In particular, for the improvement process, instead of applying a local search whose criterion is based on reaching a local optimum, a trajectory metaheuristic can be used to set a time-based stopping criterion. This option is not promising at present for the LOP with large instances because with the current amount of power, it would require setting a very short time or evolving very few generations, but as the computing power increases, this option becomes viable.
@@
In order to properly design an ILS strategy, we must select the perturbation and local search components. Given the efficiency and effectiveness of the local search based on shifts, this was left unchanged. In the case of the perturbation strategy, we analyzed several operators that have been successful in combinatorial optimization [29] and that exhibit different perturbation strengths. In particular, Scramble, Inversion, Insertion and Swap were tested. In the case of the swap strategy, we considered the version that incorporates the parameter to establish the number of 2-position swaps that are made [30]. The preliminary analyses allowed us to determine that the method based on performing multiple swaps was far superior. In particular, was enough to escape local optima, so the rest of the paper considers this operator exclusively. In fact, this operator had already shown its superiority in [28].
To extend MA-EDM (Algorithm 1) and generate MA-EDM, lines 2 and 7 are updated by replacing the local search with ILS. Algorithm 2 details each step of ILS. Each solution is intensified with local search after being perturbed, and it is accepted if it is at least as good as the best solution found. In relation to the parallel version of this algorithm (PMA-EDM), it follows the same flow presented above (see Figure 1). The only difference is that the workers execute the ILS instead of the local search.
@@ 5 Experimental Validation
The main objective of the experimental validation carried out in this research is to analyze the behavior of the leading LOP solvers in long-term executions, with the aim of understanding which design trends can help to better scale these approaches. Additionally, the new solvers designed with the most promising design decisions are applied to the most challenging LOP instances by simulating scenarios with a large amount of computation, in order to understand the room for improvement present compared to the current BKS [11], as well as to establish new levels that can be used in the future to validate new advances. The experimental analysis is conducted from three perspectives. The first focuses on identifying the best approaches for extended long-term executions. The second examines the relationship between the components of MAs and their scalability. The final utilizes HPC to analyze the performance during extended long-term executions and to establish new BKS for future comparisons.
To account for the stochastic behavior of the algorithms analyzed, the runs were repeated 30 times with distinct random seeds, and the following statistical analyses were performed to compare the results [31]. First, the Kruskal Wallis test was applied as an omnibus test. If there were any significant differences, the Mann Whitney with Hommel’s correction was used to identify the pair-wise differences. Additionally, with the aim of visualizing the results, the Nemenyi critical difference plot was used [32]. Significance levels were set to 0.05. The experimental validation was carried out in a cluster whose nodes have two Intel Xeon E5 v4 and 32 GB of RAM.
| Optimizer | Parameterization |
| ILSr | , , |
| MAr | |
| CD-RVNS | Parameter-less Algorithm |
| MA-EDM | N=200, Crossover Operator=CX |
5.1 Relative performance between population-based algorithms and trajectory algorithms
Recent best-known solutions for the LOP has been attained by considering population-based strategies. In order to confirm the superiority of population-based approaches in long-term executions, this section analyzes the behavior of four algorithms: two trajectory-based algorithms (ILSr [33] and CD-RVNS [34]) and two population-based algorithms (MAr [27] and MA-EDM [11]). These runs were conducted by setting the stopping criterion to 192 hours (eight days). Particularly, an instance of each size (, , and ) of the xLOLIB2 benchmark set was selected randomly. Table 1 shows the parameterization applied in each optimizer.
The relative performance between the strategies was similar in all the instances. Figure 2 shows the boxplots of the results for the two largest instances. The results for the smallest instances can be found in the supplementary material. MA-EDM consistently exhibits superior performance, outperforming both the other population-based algorithm and the trajectory-based algorithms and as the instances grow, the gap between the performance of MA-EDM and the rest of the strategies increase. Note that the main difference between MAr and MA-EDM is the explicit management of diversity. Thus, this experiment confirms the superiority of population-based strategies for long runs, as well as, the importance of correctly managing the diversity.
@@ 5.2 Relationship between the scalability and components of MA-EDM
In order to analyze the impact of increasing the number of generations, the performance of MA-EDM and MA-EDM was analyzed using different stopping criteria (6, 24 and 192 hours). Four instances of xLOLIB2 were chosen at random, one for each size available: 300, 500, 750 and 1000. The only parameter of MA-EDM is the population size (). Taking into account the results in [11], was used. In the case of MA-EDM, the same population size was used and the stopping criterion for ILS was set at 3.6 seconds. The preliminary analyses showed that after this run time, the improvements offered by ILS are quite limited.
Figure 3 shows the boxplots of the results obtained in the largest instances, while the results for the smallest ones can be consulted in the supplementary material. With short stopping criteria, MA-EDM is superior. This is especially clear in the 6-hour run, a case in which there are significant differences in favor of MA-EDM in every instance. However, as the computing power increases, the improvements achieved by MA-EDM are limited, and it is surpassed by MA-EDM (also in a statistically significant way). This improvement is most noticeable for larger instances. In the case of the instance of size , MA-EDM and MA-EDM evolved on average and generations, respectively. The results show that as the computing capabilities increase, expanding the number of generations is not the best design choice. In fact, in the case of the instance, despite the noticeable margin for improvement still left for MA-EDM, there were no significant differences between the results obtained in 24 hours and 192 hours.
@@
Taking into account the analyses carried out in [25], it is natural to consider MAs with large populations as an alternative to leverage the increase in computing power. Because of this, MA-EDM was run by setting the stopping criterion to 192 hours and using larger population sizes. In particular, 6 values were tested between and . The boxplots of the results obtained for the instances N-stabu75_750 and N-t70b11xx_1000 are shown in Figure 4. The results obtained by MA-EDM are also included. The statistical tests carried out confirm that in both instances, there were significant differences between the results achieved by MA-EDM and those obtained by MA-EDM, regardless of the population size used. We conclude that in this case, decreasing the number of generations through population growth is not enough to take advantage of the computing resources; instead, more drastic changes must be made by modifying some of its internal components, as is done in MA-EDM.
@@ 5.3 Analysis of performance in the very long-term
This second experiment seeks to establish new BKS and analyze whether MA-EDM scales appropriately when considering larger computation quantities than those used in the previous experiment. Since the previous experiment involved optimizations lasting up to 8 days, it is not feasible to perform this new experiment using sequential executions. Instead, PMA-EDM with 32 worker processes was used. One modification was made to the population size, consisting of establishing in order for the population size to be a multiple of the number of workers. Figure 5 shows the boxplots obtained by PMA-EDM when executed setting the stopping time between 6 hours and 26 hours for the 750 and 1000 size instances that were used in the previous experiment. Also shown are the results obtained by MA-EDM at 192 hours. The results for the 300 and 500 size instances are shown in the Supplementary Material. Note that the computation hours used by PMA-EDM range from to hours. As expected, the 6-hour parallel results were worse than the 192-hour sequential results due to the loss of performance inherent in parallel models. However, when using 10-hour runs, the median obtained by PMA-EDM was higher than that of MA-EDM. Therefore, the speedup is estimated to be between and . When considering the mean value, the efficiency is , and therefore each hour of parallel execution is estimated to correspond to hours of sequential execution. Thus, this experiment simulates situations with current computing capabilities between 153 hours and 665 hours, i.e., practically one month.
@@
In the case of the instance of size , the results continue to improve steadily, even on the longest runs. The improvement in quality obtained compared to the previous BKS is immense, thus confirming that in instances of size with the usual execution duration — at most a few days — there is considerable room for improvement. Figure 6 shows critical difference plots for these runs for the and instances. These same figures for the case of 500 are shown in the Supplementary Material of this research. In the case of 300, it is not possible to construct this plot since the same result was obtained for all the executions carried out with both algorithms. In both instances, we see that the sequential run at 192 hours ranked between the parallel runs at 6 and 10 hours, confirming the previous findings in terms of the estimated efficiency. In the 750 instance, 22-hour parallel runs were needed to significantly outperform the sequential results. When a method generates near-optimal solutions, the resources allocated have to be greatly increased to achieve significant improvements, so this, coupled with the fact that the 22-hour run ranked similar to the 26-hour run, suggests that in this instance, very high quality solutions are being reached. In the case of the instance, the sequential results at 192 hours could be significantly outperformed by all the parallel runs starting at 14 hours. In addition, the best average ranking was obtained at 26 hours, which is an indication that there is still room for improvement.
| Instance | Mean | Best | Worst | Previous BKS |
@@
| Instance | Mean | Best | Worst | BKS |
@@
6 Conclusion
The design of metaheuristics has evolved over time due to the emergence of new ideas to perform optimization processes, as well as the continuous increase in computing capabilities. This paper focuses on understanding the impact that the future increase in the availability of computational resources will have on Memetic Algorithms (MAs) in the particular case of the Linear Ordering Problem (LOP). The use of parallelizations that keep the evolutionary model intact, as well as of long-term executions, made it possible to confirm that the MA that is used today to obtain the Best-Known Solutions (BKS) in the most challenging instances of the LOP (MA-EDM), as well as other state-of-the-art optimizers, are not able to adequately leverage the increase in computational resources on a large scale. Two of the most common techniques used to take advantage of the increase in computational resources are to expand the size of the population or the number of generations. However, with a high increase in computational resources, neither of these two techniques was successful in the case of the LOP. A new heuristic optimizer (MA-EDM) was proposed based on integrating Iterated Local Search with MA-EDM, so as to avoid an excessive increase in the population size or in the number of generations, and thus increase the intensification potential. This technique was able to successfully take advantage of amounts of computational resources similar to what a modern core can execute in more than 4 months, and yielded new solutions that significantly outperform the current BKS in the 18 instances that were used in this research. These achievements confirm that as computational resources continue to increase, it will be necessary to adapt the current optimizers, and in particular it shows how promising it is to integrate MAs with techniques with great intensifying power. Accordingly, we believe that in the future there will be a significant increase in MAs that will include intensification procedures with a high computational cost. In fact, in recent years, the term matheuristics has become popular to refer to heuristics that include mathematical programming techniques in some of their components. These types of techniques were not viable in the past due to their computational cost, but the number of areas in which this type of integration has been successful is growing.
@@ 6 Conclusion
Declarations
Confict of interest The author does not have financial or non-financial interests that are directly or indirectly related to this article.